### v1

@@ 5 Simulation Results
Baselines. To validate these hypotheses, we compare Neural NOD model (7) against two baseline methods. All neural networks use the same architecture and number of parameters (fully-connected, feed-forward, and 5 hidden layers with 256 neurons each).
  • Multi-layer perceptron learned from an inverse game (MLP-IG). A mapping from feature ${x}_{t}$ to cost weights $\theta_{t}$, which is parameterized by a generic multilayer perceptron (MLP) model and learned with an inverse dynamic game. This approach is briefly explored in the prior state-of-the-art inverse game approach Liu et al. [33], which learns a tiny neural network that predicts game objectives.
  • End-to-end behavior cloning (E2E-BC). An end-to-end control policy (i.e., one that takes as input state ${x}_{t}$ and returns agents’ control ${u}_{t}$) learned with behavior cloning and supervised learning [7].
@@
  • Average accumulated steering rate difference (AASD). The accumulated steering rate difference is defined as $\sum_{k=0}^{\tilde{T}-1}|\omega_{k+1}-\omega_{k}|$ for each trial, where $\tilde{T}$ is the trial horizon, and $\omega_{k}$ is the steering rate control of the ego car at time $k$. This metric assesses the overall turning smoothness of the ego car, with a higher AASD indicating a greater likelihood of spinning out.
We will report SR and OR in terms of percentage, and AELD and AASD in terms of mean and standard deviation calculated across all simulation trials.
5.1 Synthetic dataset
In the first experiment, we learn all policies from a synthetic dataset with racing demonstrations obtained by running a game-theoretic solver. The expert game solver, inspired by ILQGame [14] with customizations to account for the Frenet coordinate [25] and track limits, computes an approximate feedback Nash equilibrium solution in real-time.
@@ 5.1 Synthetic dataset
Figure 4: Simulation snapshots and the time evolution of game cost weights, when the ego vehicle uses the baseline MLP learned from the synthetic dataset. The ego made a greedy choice by driving on the inside to stay closer to the racing line and hesitated to overtake the rival when approaching. Both factors led to a collision between the ego and the rival.
We first evaluated each policy with in-distribution rival behaviors, where the rival used an ILQGame policy whose blocking cost weight was randomized within the same range ($\theta^{r}_{\rm{bl}}\in[0,30]$) as the training data. We simulated 100 races with randomized initial states and rival cost weights, and the statistics are shown in Table 1. The proposed Neural NOD outperformed both baselines in all metrics. We then simulated another 100 randomized races against a more aggressive rival, whose blocking cost weight was randomized within the range $\theta^{r}_{\rm{bl}}\in[60,80]$. As shown in Table 2, in this more challenging, out-of-distribution setting, Neural NOD led the baselines by a significant margin in terms of safe rate, overtaking rate, and AELD. It maintained the same level of smoothness as the MLP baseline, as indicated by the tiny difference in AASD. The E2E-BC baseline yielded a significantly lower safe rate, confirming the well-known generalization issue of behavior cloned policies when deployed in previously unseen scenarios. Finally, in 100 additional randomized races where the rival used the E2E-BC policy—a shift from the game-based policy in the synthetic dataset—the Neural NOD continued to outperform the baselines in all metrics. Those results jointly validated H1. The results in Table 2 and Table 3 validated H2.
@@
Although the rival’s ILQGame cost weights were not accessible to the ego’s game policy (Neural NOD and MLP-IG), the two players’ game solvers still shared a set of common knowledge including the basis cost functions and system dynamics, which can constitute privileged information of the game-based policies over the E2E-BC baseline. Therefore, we provided additional simulation results in Table 3 with the rival using the E2E-BC policy, in which case the game-based policies no longer have an information privilege.
@@
MLP-IG $\mathbf{95\%}$ $93.68\%$ $42.87\pm 24.30$ $1.32\pm 0.33$
E2E-BC $90\%$ $88.89\%$ $39.39\pm 26.68$ $1.87\pm 0.46$
Neural NOD (ours) $\mathbf{95\%}$ $\mathbf{96.84\%}$ $\mathbf{50.85\pm 29.91}$ $\mathbf{1.29\pm 0.38}$
Table 1: Results obtained from 100 randomized races with in-distribution rival behaviors and policies trained on the synthetic dataset.
@@
MLP-IG $82\%$ $62.20\%$ $4.50\pm 37.84$ $\mathbf{1.07\pm 0.58}$
E2E-BC $66\%$ $60.61\%$ $3.64\pm 33.71$ $2.21\pm 0.47$
Neural NOD (ours) $\mathbf{91\%}$ $\mathbf{76.92\%}$ $\mathbf{7.60\pm 42.25}$ $1.09\pm 0.66$
Table 2: Results obtained from 100 randomized races in an out-of-distribution evaluation with more aggressive rival behaviors.
@@
MLP-IG $70\%$ $61.43\%$ $11.77\pm 35.00$ $1.48\pm 0.43$
E2E-BC $55\%$ $61.82\%$ $14.32\pm 35.07$ $2.52\pm 0.57$
Neural NOD (ours) $\mathbf{87\%}$ $\mathbf{74.71\%}$ $\mathbf{16.34\pm 52.49}$ $\mathbf{1.36\pm 0.36}$
Table 3: Results obtained from 100 randomized races with the rival using the behavior-cloned policy trained on the synthetic dataset.
@@
MLP-IG $78\%$ $53.85\%$ $14.49\pm 42.95$ $1.23\pm 0.64$
E2E-BC $62\%$ $75.81\%$ $12.35\pm 34.77$ $2.02\pm 0.55$
Neural NOD (ours) $\mathbf{81\%}$ $\mathbf{82.72\%}$ $\mathbf{15.94\pm 34.55}$ $\mathbf{1.09\pm 0.52}$
Table 4: Results obtained from 100 randomized races with ego policy trained on the human-generated dataset. The rival uses a game policy with randomized cost weights.
In Figure 3 and Figure 4, we examine one representative simulated race with the ego using the Neural NOD and the baseline MLP, respectively. In both cases, vehicles started from the same initial conditions, and the rival used a game policy whose cost parameters were not accessible to the ego’s policy. The ego using Neural NOD rapidly formed a strong opinion in favor of overtaking the rival from the outside. This resulted in a smooth, decisive, and safe overtaking maneuver. On the other hand, the MLP model produced a different ego behavior: the intention (encoded by $\theta^{e}_{\rm{ov}}$) to overtake the rival started high but gradually decreased as the ego approached the rival and hesitated to carry out the overtaking. Additionally, in contrast to the NOD decision, the ego greedily chose to overtake from the inside—a position closer to the time-optimal racing line but more prone to collisions due to the narrow corridor between the rival and track limit. Indeed, as the rival also attempted to defend the inside to prevent the ego from overtaking, the ego was left with no space to dodge but collided with the rival.
@@
5.2 Human-generated dataset
Next, we learn all policies from a dataset with racing demonstrations performed by human drivers. To obtain such racing data, we leveraged a driving simulator (shown in the upper right corner of Figure 2), where the ego car was driven by human participants of various skill levels against a rival using a reactive RL-based policy in Carla [11]. For evaluation, we simulated 100 races with randomized initial states and a rival using a game policy with randomized cost weights inaccessible to the ego’s planner. The racing statistics are shown in Table 4. The Neural NOD again outperformed both baselines in all metrics. This result coincides with those obtained based on the synthetic dataset, confirming the ability of our proposed inverse game framework to learn a Neural NOD from noisy human data (H3). In Figure 5, we compare a simulation trajectory generated by Neural NOD with the groundtruth replay. Under this initial condition when the rival tried to defend from the inside (left), the Neural NOD decided to attack from the outside (right), controlling the ego car to successfully overtake the rival. Note that this decision closely resembles the groundtruth human demonstration, also displayed in Figure 5.
5.3 The full endurance race
@@
6 Limitations and Future Work
In this paper, we consider a restricted setting where Neural NOD is used to automatically tune the (continuous) cost weights of a dynamic game solver. As shown in recent work [22], an opinion state can be interpreted as a simplex (e.g., a probability vector), and thus a Neural NOD model is also capable of coordinating agents over integer-valued options such as leadership assignment in a Stackelberg setting [26, 19]. In addition, while we have demonstrated that the attention $\lambda$ governed by the black-box DNN generally performed well for car racing, we may use an explicit attention model that enables excitable decision-making by adding an extra slower negative feedback loop [28, Sec. 7.2.2]. Such enhanced flexibility in excitable decision-making allows agents to timely forget prior decisions and thus never get stuck in a decision that may no longer be safe and/or efficient as the environment has evolved rapidly. Finally, we see an open opportunity to use Neural NOD for shared autonomy, e.g., AI-assisted car racing [49, 38], in which the human and robot simultaneously provide control inputs to the system while interacting with other agents. By modeling the intents of the human and robot as opinion states and planning robot motion in the joint intent–physical space [18], we can achieve not only value alignment, i.e., the robot infers and adopts the human’s goals, but also automation transparency, i.e., the human is aware of the robot’s current intent, both rapidly and decisively with Neural NOD.
7 Conclusion
In this paper, we introduced a Neural NOD model for game-theoretic robot motion planning in split-second, and an inverse game approach to learn such a model from data. We also provided a constructive procedure to adjust Neural NOD parameters online such that breaking of indecision is guaranteed. Through extensive simulation studies of autonomous racing based on real-world circuit and human-generated interaction data, we demonstrated that a dynamic game policy guided by a Neural NOD can consistently outperform state-of-the-art imitation learning and data-driven inverse game policies. Beyond fully-automated robot interactions, the framework of Neural NOD provides an open opportunity to enable fast, decisive, and transparent value alignment for broader human-AI shared control systems.
Acknowledgements
This work is supported by Toyota Research Institute (TRI). It solely reflects the opinions and conclusions of its authors and not TRI, or any other Toyota entity. The authors thank Jingqi Li and Xinjie Liu for very helpful discussions on inverse dynamic games, and Alessio Franci for generously providing insights into nonlinear opinion dynamics.

### v2

@@ IV Simulation Results
Baselines. We compare our approach with two baseline methods. All neural networks have five fully-connected layers with 256 neurons each. The network features include all vehicles’ physical states.
  • Multi-layer perceptron learned from an inverse game (MLP-IG). A mapping from feature ${x}_{t}$ to cost weights $\theta_{t}$, parameterized by a multilayer perceptron (MLP) model and learned with an inverse dynamic game. This approach is briefly explored in the prior state-of-the-art inverse game approach [liu2023learning], which learns a DNN that predicts game objectives.
  • End-to-end behavior cloning (E2E-BC). An end-to-end control policy (i.e., one that takes as input state ${x}_{t}$ and returns agents’ control ${u}_{t}$) learned with behavior cloning and supervised learning [bojarski2016end].
@@
We will report SR and OR in percentage, and AELD with mean and standard deviation calculated across all trials.
IV-A Synthetic dataset
In the first experiment, we learn all policies from a synthetic dataset with racing demonstrations obtained with a model-based trajectory optimizer. We first evaluated each policy with in-distribution rival behaviors, where the rival used an ILQGame policy whose blocking cost weight was randomized within the same range ($\theta^{r}_{\rm{bl}}\in[0,30]$) as the training data. We simulated 100 races with randomized initial states and rival cost weights, and the statistics are shown in Table I. The proposed Neural NOD outperformed both baselines in all metrics.
@@ IV-A Synthetic dataset
We then simulated another 100 randomized races against a more aggressive rival, whose blocking cost weight was randomized within the range $\theta^{r}_{\rm{bl}}\in[60,80]$. As shown in Table II, in this more challenging, out-of-distribution setting, Neural NOD led the baselines by a significant margin in terms of safe rate, overtaking rate, and AELD. The E2E-BC baseline yielded a significantly lower safe rate, confirming the well-known generalization issue of behavior cloned policies when deployed in previously unseen scenarios.
Finally, in 100 additional randomized races where the rival used the E2E-BC policy, the Neural NOD continued to outperform the baselines in all metrics. Those results jointly validated H1. Results in Table II and Table III validated H2.
@@
Although rival’s cost weights were not accessible to the ego, the solvers still shared a set of common knowledge (e.g., the basis cost functions), which can constitute privileged information of the game-based policies over E2E-BC. In Table III, we provide results with the rival using the E2E-BC policy, in which case the game-based policies no longer have an information privilege.
@@
MLP-IG $\mathbf{95\%}$ $93.68\%$ $42.87\pm 24.30$
E2E-BC $90\%$ $88.89\%$ $39.39\pm 26.68$
Neural NOD (ours) $\mathbf{95\%}$ $\mathbf{96.84\%}$ $\mathbf{50.85\pm 29.91}$
TABLE I: Results obtained from 100 randomized races with in-distribution rival behaviors and policies trained with synthetic data.
@@
MLP-IG $82\%$ $62.20\%$ $4.50\pm 37.84$
E2E-BC $66\%$ $60.61\%$ $3.64\pm 33.71$
Neural NOD (ours) $\mathbf{91\%}$ $\mathbf{76.92\%}$ $\mathbf{7.60\pm 42.25}$
TABLE II: Results obtained from 100 randomized races in an out-of-distribution evaluation with more aggressive rival behaviors.
@@
MLP-IG $70\%$ $61.43\%$ $11.77\pm 35.00$
E2E-BC $55\%$ $61.82\%$ $14.32\pm 35.07$
Neural NOD (ours) $\mathbf{87\%}$ $\mathbf{74.71\%}$ $\mathbf{16.34\pm 52.49}$
TABLE III: Results obtained from 100 randomized races with the rival using the behavior-cloned policy trained with synthetic data.
@@
MLP-IG $78\%$ $53.85\%$ $14.49\pm 42.95$
E2E-BC $62\%$ $75.81\%$ $12.35\pm 34.77$
Neural NOD (ours) $\mathbf{81\%}$ $\mathbf{82.72\%}$ $\mathbf{15.94\pm 34.55}$
TABLE IV: Results obtained from 100 randomized races with ego policy trained on the human-generated dataset. The rival uses a game policy with randomized cost weights.
@@
In Figure 3, we examine one representative simulated race with the ego using the Neural NOD. The vehicles started from the same initial conditions, and the rival used a game policy whose cost parameters were not accessible to the ego’s policy. The ego agent rapidly formed a strong opinion in favor of overtaking the rival from the outside. This resulted in a smooth, decisive, and safe overtaking maneuver.
IV-B Human-generated dataset
Next, we learn all policies from a dataset with racing demonstrations performed by human drivers. To obtain such racing data, we leveraged a driving simulator (shown in the upper right corner of Figure 2), where the ego car was driven by human participants of various skill levels against a rival using a reactive RL-based policy in CARLA [dosovitskiy2017carla]. For evaluation, we simulated 100 races with randomized initial states and a rival using a game policy with randomized cost weights inaccessible to the ego’s planner. The racing statistics are shown in Table IV. The Neural NOD again outperformed both baselines in all metrics. This result coincides with those obtained based on the synthetic dataset, confirming the ability of our proposed inverse game framework to learn a Neural NOD from noisy human data (H3). In Figure 4, we compare a simulation trajectory generated by Neural NOD with the groundtruth replay. Under this initial condition, when the rival tried to defend from the inside (left), the Neural NOD decided to attack from the outside (right), controlling the ego car to successfully overtake the rival. Note that this decision closely resembles the groundtruth human demonstration, also displayed in Figure 4.
IV-C The full endurance race
@@
V Limitations and Future Work
In this paper, we consider a restricted setting where Neural NOD is used to automatically tune the (continuous) game cost weights. As shown in recent work [hu2023emergent], an opinion state can be interpreted as a simplex. Therefore, a Neural NOD model is also capable of coordinating agents over integer-valued options such as leadership assignment in a Stackelberg setting [khan2023leadership, hu2024plays]. In addition, while attention $\lambda$ governed by the black-box DNN generally performs well for car racing, we may use an explicit attention model to enable excitable decision-making by, for example, adding an extra slower negative feedback loop [leonard2024fast, Sec. 7.2.2]. Such enhanced flexibility would allow agents to forget prior decisions to prevent sticking to decisions that may no longer be safe and/or competitive as the environment evolves rapidly. Finally, we see an open opportunity to use Neural NOD for shared autonomy, e.g., AI-assisted car racing [TANIGUCHI2014, decastrodreaming], in which the human and robot simultaneously provide control inputs to the system while interacting with other agents. By modeling the agents’ intents as opinion states and planning robot motion in the joint intent–physical space [hu2024doxo], we can achieve not only value alignment, i.e., the robot infers and adopts the human’s goals, but also automation transparency, i.e., the human is aware of the robot’s current intent, both rapidly and decisively with Neural NOD.
VI Conclusion
In this paper, we introduced a Neural NOD model for game-theoretic robot motion planning in a split-second, and an inverse game approach to learn such a model from data. We also provided a constructive procedure to adjust Neural NOD parameters online such that breaking of indecision is guaranteed. Through extensive simulation studies of autonomous racing based on real-world circuit and human-generated interaction data, we demonstrated that a dynamic game policy guided by a Neural NOD can consistently outperform state-of-the-art imitation learning and data-driven inverse game policies.
[…]